# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 4.3 Kac-scale regularity of mesoscopic block averages
$\left\|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)\right\|_{L^{\infty}}\leq\varepsilon_{\gamma}\left\|\nabla_{\varepsilon_{\gamma}}\phi\right\|_{L^{\infty}}\leq C_{\phi}\,\varepsilon_{\gamma}.$ (29)
By estimate 27, we have
$\sup_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq\left\|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\right\|_{L^{\infty}}\leq C\,\frac{\gamma}{\delta_{\gamma}\varepsilon_{\gamma}}.$ (30)
Since $\kappa_{\gamma}(0)=\gamma\mathfrak{K}(0)>0$,
$\kappa_{\gamma}(0)\,|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|^{2}\leq\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}.$
Thus
$|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|\leq\frac{1}{\sqrt{\gamma\mathfrak{K}(0)}}\left(\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}.$
Then by Jensen’s inequality,
$\mathcal{E}_{\gamma}^{(1)}\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\frac{\gamma}{\delta_{\gamma}}\frac{1}{\sqrt{\gamma\mathfrak{K}(0)}}\left(\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}$
$\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\,\frac{1}{N}\sum_{i\in\Lambda_{N}}\left(\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}$
$\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\left[\frac{1}{N}\sum_{i}\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right]^{1/2}.$ (31)
Hence, by Lemma 11 with $J\equiv 1$, we have $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{E}_{\gamma}^{(1)}]=0.$
We next treat $\mathcal{E}_{\gamma}^{(2)}$. By the deterministic kernel matching estimate,
@@
$\mathcal{E}_{\gamma}^{(2)}\leq C_{\phi}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}N\frac{\gamma}{\delta_{\gamma}}(\gamma L_{\gamma})=C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}.$
Since $\frac{\varepsilon_{\gamma}^{2}\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}=\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\,\gamma^{3/2}L_{\gamma}$ and $\gamma L_{\gamma}\to 0$, we have $\gamma^{3/2}L_{\gamma}\to 0$, we have $\lim_{\gamma\downarrow 0}\mathcal{E}_{\gamma}^{(2)}=0.$ Combining with the convergences of $\mathcal{E}_{\gamma}^{(1)}$ and $\mathcal{E}_{\gamma}^{(2)}$, we obtain
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl[(m_{i}^{L_{\gamma}}(\sigma(t)))^{2}-\delta_{\gamma}^{2}X_{\gamma}(t,\varepsilon_{\gamma}i)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|=0.$ (32)
It remains to control $r_{\gamma}^{(2)}$. Since $|\Lambda_{N}|=N=\varepsilon_{\gamma}^{-1}$, estimates (30) yields that
$\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq C_{\phi}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\,N\,\frac{\gamma}{\delta_{\gamma}}=C_{\phi}\,\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}.$
Using the uniform bound $\sup_{\gamma\downarrow 0}\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}<\infty$ and $\sup_{|m|\leq 1}|\rho_{\gamma,L_{\gamma}}(m)|\leq C_{\beta}\gamma^{2}L_{\gamma}^{3}$, we obtain
$\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}|r_{\gamma}^{(2)}(\phi;t)|\,dt\right]\leq C_{\phi,\beta,T}\,\gamma^{2}L_{\gamma}^{3}\sup_{\gamma\downarrow 0}\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\to 0,\quad\text{as}\quad\gamma^{2}L_{\gamma}^{3}\to 0.$
@@
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=$ $\langle X_{\gamma}(0),\phi\rangle_{\gamma}-A_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds$
@@
where $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\Bigl[\sup_{t\leq T}|R_{\gamma}(t,\phi)|\Bigr]=0.$, $\nu_{\gamma\to\nu}$ $A_{\gamma}:=A^{\prime}_{\gamma}+A^{\prime\prime}_{\gamma}\to A=\lambda\frac{(1+\beta)}{2}$, and $\lambda_{\gamma}\to\lambda$. Equivalently,
@@ 5 Convergence of the noise term
In this section we analyze the limit of noise term in the one-dimensional Ising–Kac–Kawasaki dynamics. Then we show that the associated discrete stochastic convolution term $Z_{\gamma}$ converge to the stochastic convolution term $Z(t):=\sqrt{\sigma_{*}}\int_{0}^{t}S(t-r)\nabla\cdot\xi(r)dr.$
@@ 5.1 Convergence of the discrete martingale
Recall that the associated Dynkin martingale is
$M_{t}^{\gamma}(\phi):=\langle X_{\gamma}(t),\phi\rangle_{\gamma}-\langle X_{\gamma}(0),\phi\rangle_{\gamma}-\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma(s/\alpha_{\gamma}))\,ds.$
where $\phi$ is the smooth test function in $C^{\infty}(\mathbb{T})$. The structure is parallel to the multiscale analysis of the drift part, but simpler: only the zeroth-order local equilibrium coefficient survives in the quadratic variation.
@@
Using the induced increment of the Kac-smoothed field (12), the jump of the field pairing is
$\delta_{i,i+1}^{\gamma}\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}$ $:=\langle X_{\gamma}(\sigma^{i,i+1}),\phi\rangle_{\gamma}-\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}\sum_{k\in\Lambda_{N}}d_{x}\Bigl(\kappa_{\gamma}(k-i-1)-\kappa_{\gamma}(k-i)\Bigr)\phi(\varepsilon_{\gamma}k)$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}\,d_{i}\Bigl[(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}(i+1))-(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Bigr]$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}^{2}\,d_{i}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
The pure-jump martingale $M^{\gamma}(\phi)$ has quadratic variation is
$\langle M^{\gamma}(\phi)\rangle_{t}=$ $\int_{0}^{t}\alpha_{\gamma}^{-1}\sum_{i\in\Lambda_{N}}c_{i}\Big(\sigma\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\bigl(\delta_{i,,i+1}^{\gamma}\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}\bigr)^{2}\,ds$
$=$ $\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}d^{2}_{i}\big(\frac{s}{\alpha_{\gamma}}\big)c_{i}\Big(\sigma\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\Bigl[\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Bigr]^{2}\,ds,$
Similar with the second-order Boltzmann–Gibbs principle for $d^{2}_{i}$, we now consider $c_{i}(\sigma)d^{2}_{i}$.
A direct computation shows that
@@
$\sup_{i,\sigma}|\Delta_{i}H_{\gamma}(\sigma)|\leq C\gamma.$
@@
the logistic map is globally Lipschitz. Therefore
$c_{i}(\sigma)=F(\Delta_{i}H_{\gamma}(\sigma))=\frac{1}{2}+r_{i}^{\gamma}(\sigma),\qquad\sup_{i,\sigma}|r_{i}^{\gamma}(\sigma)|\leq C\gamma.$
@@
Note that $\widetilde{r}_{i}^{\gamma}$ is not a strictly local observable; it will be treated as a small uniform error term.
@@
$\quad+\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma\!\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds.$
Then by estimate (29), we have
$\Big|\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma\!\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds\Big|\leq C_{\phi}\,t\,\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\quad\text{as }\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\to 0.$
By Theorem 5 and Lemma 10, the local observable $d_{i}^{2}$ may be replaced by the local equilibrium average
$\Phi_{d_{0}^{2},L_{\gamma}}^{\gamma}(m)=a_{0,\gamma}+a_{2,\gamma}m^{2}+\rho_{\gamma,L_{\gamma}}(m).$
Therefore
$\langle M^{\gamma}(\phi)\rangle_{t}=\mathcal{Q}_{\gamma}^{(0)}(\phi;t)+\mathcal{Q}_{\gamma}^{(2)}(\phi;t)+\mathcal{R}_{\gamma}^{(\rho)}(\phi;t)+o_{\gamma}(1),$
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t):=\widetilde{a}_{0,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t):=\widetilde{a}_{2,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\bigl(m_{i}^{L_{\gamma}}(\sigma(s/\alpha_{\gamma}))\bigr)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
@@
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t)=\widetilde{a}_{2,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}X_{\gamma}(s,\varepsilon_{\gamma}i)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}ds+o_{\gamma}(1).$
@@
Then by $L^{2}$ bound (28) for $X_{\gamma}$ and $\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\to 0,$ we conclude that
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,\widetilde{a}_{0,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}.$
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,\left(\widetilde{a}_{0,\gamma}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\right)\left(\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|^{2}\right)\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx.$

### 05 Apr 2026 v2

@@ 4.3 Kac-scale regularity of mesoscopic block averages
$\left\|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)\right\|_{L^{\infty}}\leq\left\|\nabla_{\varepsilon_{\gamma}}\phi\right\|_{L^{\infty}}\leq C_{\phi}.$ (36)
By estimate (18), we have
$\sup_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq C_{\phi}\left\|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\right\|_{L^{\infty}}\leq C\ \frac{\gamma}{\delta_{\gamma}\varepsilon_{\gamma}}.$ (37)
Then by Lemma 13, we have $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{E}_{\gamma}^{(1)}]=0.$
We next treat $\mathcal{E}_{\gamma}^{(2)}$. By the deterministic kernel matching estimate, the Kac kernel convolved with block-to-site error
@@
$\mathcal{E}_{\gamma}^{(2)}\leq C_{\phi}\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}}N\frac{\gamma^{2}L_{\gamma}}{\delta_{\gamma}}=C_{\phi}\frac{\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\to 0\quad\text{as}\ \gamma\to 0.$
Combining with the convergences of $\mathcal{E}_{\gamma}^{(1)}$ and $\mathcal{E}_{\gamma}^{(2)}$, we obtain
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl[(m_{i}^{L_{\gamma}}(\sigma\big(t/\alpha_{\gamma}\big)))^{2}-\delta_{\gamma}^{2}X_{\gamma}(t,\varepsilon_{\gamma}i)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|=0.$ (38)
It remains to control $r_{\gamma}^{(2)}$. Since $|\Lambda_{N}|=2N+1=\varepsilon_{\gamma}^{-1}$, estimates (37) yields that
$\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq C_{\phi}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\,N\,\frac{\gamma}{\varepsilon_{\gamma}\delta_{\gamma}}=C_{\phi}\,\frac{\gamma}{\alpha_{\gamma}\delta_{\gamma}}.$
Using the bound $\sup_{|m|\leq 1}|\rho_{\gamma,L_{\gamma}}(m)|\leq C_{\beta}\gamma^{2}L_{\gamma}^{3}$, we obtain
$\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}|r_{\gamma}^{(2)}(\phi;t)|\,dt\right]\leq C_{\phi,\beta,T}\sup_{\gamma\downarrow 0}\frac{\gamma^{3}L_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}}\to 0,\quad\text{as}\quad\gamma^{2}L_{\gamma}^{3}\to 0.$
@@
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=$ $\langle X_{\gamma}(0),\phi\rangle_{\gamma}-\big(A^{\prime}_{\gamma}+A^{\prime\prime}_{\gamma}\big)\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds$
@@
where $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\Bigl[\sup_{t\leq T}|R_{\gamma}(t,\phi)|\Bigr]=0.$, $\nu_{\gamma\to\nu}$ $A^{\prime}_{\gamma}+A^{\prime\prime}_{\gamma}:=A_{\gamma}\to A=\lambda\frac{(1+\beta)}{2}$, and $\lambda_{\gamma}\to\lambda$. Equivalently,
@@ 5 Convergence of the noise term
In this section we identify the martingale part in the Dynkin decomposition and show that, after the conservative replacement of the local mobility, it converges to the Gaussian noise term in the limiting stochastic Cahn–Hilliard equation. More precisely, for every smooth test function $\phi\in C^{\infty}(\mathbb{T})$, we study the Dynkin martingale
$M_{t}^{\gamma}(\phi):=\langle X_{\gamma}(t),\phi\rangle_{\gamma}-\langle X_{\gamma}(0),\phi\rangle_{\gamma}-\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)\Bigl(\sigma^{N}(\alpha_{\gamma}^{-1}s)\Bigr)\,ds.$
Then by using the convergence of semigroup, we will show that the associated discrete stochastic convolution $Z_{\gamma}:=\int^{t}_{0}S_{\gamma}(t-s)M_{\gamma}(t)$ converge to the stochastic convolution term $Z(t):=\sigma_{*}\int_{0}^{t}S(t-r)\nabla\cdot\xi(r)dr.$
@@ 5.1 Convergence of the discrete martingale
The goal of this subsection is twofold:
  1. (i)
    to compute the predictable quadratic variation of $M^{\gamma}(\phi)$ and identify its limit as $\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\nabla\phi(x)|^{2}\,dx;$
  2. (ii)
    to deduce, by a martingale central limit theorem, that the family $\{M^{\gamma}(\phi)\}_{\gamma}$ converges to the centered Gaussian martingale associated with the conservative noise $\sigma_{*}\nabla\xi$.
The structure in the Dynkin martingale $M_{t}^{\gamma}(\phi)$ is parallel to the multiscale analysis of the drift part, but simpler: only the zeroth-order local equilibrium coefficient survives in the quadratic variation.
@@
We first compute the predictable quadratic variation. At a Kawasaki jump across the bond $(i,i+1)$, the configuration changes from $\sigma$ to $\sigma^{i,i+1}$, and (14) yields
$h_{\gamma}(\sigma^{i,i+1};k)-h_{\gamma}(\sigma;k)=d_{i}(\sigma)\Bigl[\kappa_{\gamma}(k-i-1)-\kappa_{\gamma}(k-i)\Bigr].$
Therefore
$\delta_{i}\langle X_{\gamma},\phi\rangle_{\gamma}(\sigma)$ $:=\langle X_{\gamma}(\sigma^{i,i+1}),\phi\rangle_{\gamma}-\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}$
$=\frac{\varepsilon_{\gamma}}{\delta_{\gamma}}\sum_{k\in\Lambda_{N}}\Bigl(h_{\gamma}(\sigma^{i,i+1};k)-h_{\gamma}(\sigma;k)\Bigr)\phi(\varepsilon_{\gamma}k)$
$=-\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\,d_{i}(\sigma)\,\nabla_{\varepsilon_{\gamma}}\!\bigl(\mathcal{K}_{\gamma}^{\varepsilon}\phi\bigr)(\varepsilon_{\gamma}i).$
Hence, by the standard bracket formula for pure-jump Dynkin martingales,
$\langle M^{\gamma}(\phi)\rangle_{t}=\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}d^{2}_{i}\big(\frac{s}{\alpha_{\gamma}}\big)c_{i}\Big(\sigma^{N}\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\Bigl[\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Bigr]^{2}\,ds.$ (39)
We set
$d_{i}(\sigma):=\sigma_{i}-\sigma_{i+1},\quad q_{i}(\sigma):=d_{i}(\sigma)^{2}\,c_{i}(\sigma),\quad i\in\Lambda_{N},$
where
$c_{i}(\sigma)=F\!\bigl(\beta\,\Delta_{i}H_{\gamma}(\sigma)\bigr),\qquad F(z):=\frac{1}{1+e^{z}}.$
The quantity $q_{i}$ is the local mobility observable entering the predictable quadratic variation of the martingale. Similar with the second-order Boltzmann–Gibbs principle for $d^{2}_{i}$, we now consider
$q_{i}(\sigma):=d_{i}(\sigma)^{2}\,c_{i}(\sigma)=d_{i}(\sigma)^{2}F\!\bigl(\beta\,\Delta_{i}H_{\gamma}(\sigma)\bigr).$
We next separate the mobility into a leading symmetric part and a small remainder. A direct computation shows that
@@
$\sup_{i,\sigma}|\beta\Delta_{i}H_{\gamma}(\sigma)|\leq C\gamma.$
@@
the logistic map is globally Lipschitz.
The Taylor expansion of $F$ at the origin gives
$c_{i}(\sigma)=F(\beta\Delta_{i}H_{\gamma}(\sigma))=\frac{1}{2}+r_{i}^{\gamma}(\sigma),\qquad\sup_{i,\sigma}|r_{i}^{\gamma}(\sigma)|\leq C\gamma.$
@@
Note that $\widetilde{r}_{i}^{\gamma}$ is not a strictly local observable; it can be treated as a small uniform error term.
@@
$\quad+\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma^{N}\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds$
$=$ $\frac{1}{2}\mathcal{Q}_{\gamma}(\phi;t)+\mathcal{R}_{\gamma}^{\mathrm{mob}}(\phi;t).$ (40)
We first show remainder term $\mathcal{R}_{\gamma}^{\mathrm{mob}}(\phi;t)$ is negligible. By estimate (36), as $\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\to 0$ we have
$\big|\mathcal{R}_{\gamma}^{\mathrm{mob}}(\phi;t)\big|=\Big|\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma^{N}\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds\Big|\leq C_{\phi}\,t\,\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}.$
By Theorem 6 and Lemma 12, $d_{i}^{2}$ can be replaced by the local equilibrium average at the mesoscopic block scale $L_{\gamma}$ with $L_{\gamma}/N\to 0$ as $\gamma\to 0$. More precisely,
$\Phi_{d_{0}^{2},L_{\gamma}}^{\gamma}(m)=a_{0,L,\gamma}+a_{2,L,\gamma}m^{2}+\rho_{\gamma,L}(m)$
in weak sense. Therefore
$\mathcal{Q}_{\gamma}(\phi;t)=\mathcal{Q}_{\gamma}^{(0)}(\phi;t)+\mathcal{Q}_{\gamma}^{(2)}(\phi;t)+\mathcal{R}_{\gamma}^{(\rho)}(\phi;t),$
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t):=a_{0,L,\gamma}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t):=a_{2,L,\gamma}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\bigl(m_{i}^{L_{\gamma}}(\sigma(s/\alpha_{\gamma}))\bigr)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
@@
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t)=a_{2,L,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}X_{\gamma}(s,\varepsilon_{\gamma}i)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}ds+o_{\gamma}(1).$
@@
By bound 18, we estimate $\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}X_{\gamma}(s,\varepsilon_{\gamma}i)^{2}\leq C\frac{\gamma^{2}}{\delta_{\gamma}^{2}}$. Then by $\frac{\gamma^{2}\varepsilon_{\gamma}^{3}}{\delta^{2}_{\gamma}\alpha_{\gamma}}\to 0,$ we conclude that
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,a_{0,L,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}.$
@@
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,\left(a_{0,L,\gamma}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\right)\left(\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|^{2}\right)\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx.$